Reports 1-1 of 1 Clear search Modify search
AdV-ISC (Commissioning up to first full interferometer lock)
bersanetti, maggiore, amagazzu - 1:26 Thursday 13 May 2021 (51771) Print this report
ISC shift: CARM/DARM handoff to B7/B8

The goal of the shift was to study the handoff of CARM/DARM from the beating signals to the B7/B8 DC recombinations while still having around 300 Hz of CARM offset.

After the work in the Laser Lab ended, there was the need to restart the green at NEB, and then another realignment of the ALS WARM in CEB was needed. After that we could recover the lock acquisition (a tuning of the lock thresholds for both North and West ALS was needed) and we could start the activity, around 16:00 UTC:

  • We relocked the arms in the CARM/DARM basis on the beating signals on resonance for the IR at 16:20 UTC; one minute later, a + 3 kHz offset was added to CARM;
  • at this moment we had some short instability of the INJ system, with a couple of unlocks and the MC transmitted power going down for a short time; the system then recovered itself;
  • we went back to the same configuration again at 16:37 UTC (when we unlocked) and then at 16:47 UTC;
  • we started to work on the DRMI which, after a first pre-alignment of both the SR and PR mirrors, could lock rather easily and in a reproducible way for the entire shift, without any change in loop parameter, up to the 3f handoff of both the LSC and ASC loops; after a short unsuccessfull test, the BS drift control was left open for the rest of the shift;
  • once starting to reduce the CARM offset, the only change of gain which was needed was for PRCL, from 0.15 to 0.06; MICH could benefit from a higher gain (UGF ~ 12 Hz) but doing so increased the unlocks due to saturations, so it was tested only twice and the loop was kept with such UGF in following trials;
  • for the most part of the shift, we could reliably move from 3 kHz CARM offset to 300 Hz in a single 10 s long ramp, while decreasing the PRCL gain and blocking the B1p and B4 triggers at the same time;
  • then we started to study the handoff of CARM and DARM to respectively LSC_B7_B8_DC_SUM_SQRT and LSC_B7_B8_DC_DIF_NORM; while doing this we suffered from an unlock at around 18:24 UTC; the cause is not clear, but we observed the NI_F7 and NI_PAY flags on the DMS becoming grey just after this event;
  • we continued this study, and we made the handoff of DARM at 19:10:10 UTC, using the DARM UGF line and recent measurements for the sensing and offset changes; we unlocked while doing the same for CARM;
  • we made another handoff of DARM at 19:47:50 UTC, and we unlocked shortly after, maybe because of the big 150 Hz noise present on the B7/B8 reconstructed signals; there is the already seen noise bump in the 100-200 Hz region, but the highest noise was at 150 Hz sharp;
  • at around 20:18:50 UTC we did another handoff, this time of CARM; then at 20:23:05 UTC we also made the DARM one, immediately reducing the line amplitude to decrease the noise; again, we unlocked shortly after; about this handoff:
    • Figure 1 shows the timeline;
    • Figure 2 shows the spectra of the Err/Corr signals when both CARM/DARM where on the beating signals (purple traces) and on the B7/B8 reconstructed signals (blue traces); the 87.1 Hz line in DARM is the UGF line we injected;
    • Figure 3 shows the DARM spectrum in four different cases: BLACK: DARM on beating signal, CARM on beating signal; PURPLE: DARM on B7/B8, CARM on beating signal (bump and 150 Hz noise as described above); GREEN: DARM on beating signal, CARM already on B7/B8; BLUE: DARM on B7/B8, CARM already on B7/B8;
    • in all the trials, the DARM UGF remained constant, around 14-15 Hz;
    • the parameters for the change of calibration and offset could vary by a 10-20% for each different trial, which could be related to the uncertainty we have on the determination of the IR resonance with the beating signals; it could be worth to check and calibrate those reconstructed signals while still locked on the IR, in a similar way to what we do with the beating signals themselves; this would require putting some offsets on the single arms lock in order to move away a bit from the IR resonance, so we need to check if it is feasible at all.

During the latest part of the shift the INJ system started again to have troubles, with frequent, slow dips on the PMC power, followed by some PSTAB instability; then the system recovered just as slowly, or immediately in case the IMC unlocked.

While doing this I noticed the ALS reflected powers (ALS_*EB_PD_REFL_LP) being quite higher than usual and following a strange trend. In Figure 4 there is such trend: the first jumps on the BEAT_DC signals are due to the lock/unlock of the ALS, but the trend is visible. The trend of the PMC power is opposite to the one of the ALS_CEB_PD_GREEN_MONI signal (zoom in Figure 5).

In agreement with the INJ crew, we left all the locking nodes in DOWN, and also INJ_MAIN in DOWN and paused.

Images attached to this report
Search Help